please help me Interpret the MS data

16 26 27 m/z Rel Abun 28 2.4 29 38 38 1.4 39 1. 57 57 Abun m/z Rel Abun 58 4.4 5971 0.08 0.04 72 0.01 2.8 100
Answer #1

The m/z of the required molecule = 72 (molecular ion peak)

There is a methyl fragment, which is indicated by m/z= 57 (base peak)

72 - 57 = 15, that corresponds to CH3 fragment.

There is a peak at m/z = 29, which corresponds to ethyl fragment (CH3-CH2)

57 - 29 = 28, which corresponds to carbonyl fragment (C=O)

Therefore, the structure of the molecule can be drawn as follows.

CH3-CH2-C(=O)-CH3 (Butan-2-one)